ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

3. Number Theory and Arithmetic Algebraic Geometry

1991 MS Classification: 11 Lanchon, Pierre, SNECMA, France Similarities in digit sequences in π and e/Correspondances entre des séquences de chiffres de π et e Coincidences exist between certain of the digit sequences of universal numbers. Are the corresponding approaching likelinesses due to pure chance or do they stem from an intelligible structure? Exemples: in binary numeration, π bits from the 7^{th} to the 25^{th}, read from right to left, correspond to the first 19 bits of e. For the function Ahmespb50(x) = ((x + 1)/x)^{(x+1)} whose limit is e when x tends towards infiniy, the fixed point is 3.1410 … = π/1.00017 . . .. The fixed point of the function Rhindpb50(x) = (1 + 1/4 + 1/4^{2} + 1/4^{3} + 1/4^{(x+1)})^{4} is 3.1415898 … = π/1.000 000 88 . . .. A natural parametrisation of the graph of this function leads to a point of abscissa x = 3.14159294 … = 1.000 000 093 … x π. Il existe des coïncidences de certaines séquences de chiffres de nombres universels. Les identités approchées correspondantes sont-elles attribuables au hasard, ou sont-elles le signe d’une structure intelligible? Exemples: en binaire, les bits 7 à 25 de π, lus de droite à gauche, correspondent aux 19 premiers bits de e. La fonction Ahmespb50(x) = ((x+1)/x)^{(x+1)} a pour limite e quand x tend vers l’infini, et pour point fixe 3.1410 … = π/1.00017 . . .. La fonction Rhindpb50(x) = (1 + 1/4 + 1/4^{2} + 1/4^{3} + 1/4^{(x+1)})^{4} a pour point fixe 3.1415898 … = π/1.000 000 88 . . .. Un paramétrage naturel de son graphe conduit à un point remarquable d’abscisse x = 3.14159294 … = 1.000 000 093 … x π.